# LAW 2025: Bridging Language, Agent, and World Models

### v3

@@ 4.2 Decomposing by Positive and Negative Responses
In this formulation, the terms $A_{q,i}^{+}$ and $A_{q,i}^{-}$ serve as adaptive weights for the positive and negative components of the objective, respectively. These relative advantage scores implicitly encode the difficulty of the question $q$. When a question is either very easy or very hard for the model, the variance among the responses tends to be low, leading GRPO to assign smaller weights to both the positive and negative terms. In contrast, for questions of moderate difficulty, where response variance is higher, GRPO assigns proportionally larger weights to both terms, thus emphasizing learning from such examples.
@@ 5 Length Bias in GRPO
This length bias arises from structural assumptions in the learning framework, specifically, dividing the advantage uniformly across tokens and scaling it with response length. We empirically show that this bias also appears in F-ISFT+-, similar to GRPO. Notably, F-ISFT+-, which is trained on both correct and incorrect responses, produces longer outputs than F-ISFT+, which is trained only on positive samples. Yet, the longer responses are often interpreted as signs of ’learning to reason’ (guo2025deepseek), even though they arise from disproportionately longer incorrect responses caused by structural assumptions in the degenerate LLM-MDP framework. We also provide further analysis on length bias in the Appendix A.
@@ 6.1 Experimental Setup
Datasets: We present our analysis on two datasets - 1) the GSM8K dataset (cobbe2021gsm8k), which is a widely used benchmark dataset consisting of grade school math problems, designed to evaluate the reasoning capabilities of large language models. It contains 8.5K problems, each paired with a question and an answer. The dataset is divided into 7.5K training problems and 1K test problems. And, 2) Countdown dataset (wikipediaCountdowngame), which is a generalized version of the classic 24 Game (yang2022chain), where the objective is to combine a set of input numbers using basic arithmetic operations (+, –, ×, ÷) to reach a specified target number. In this dataset, each problem consists of 3 to 4 two-digit input numbers, with the target number also being a two-digit number. The dataset contains 9K examples, split into 8K training instances and 1K test instances. We also report the accuracy on the MATH-500 dataset, a 500-problem test set evaluated using models trained on GSM8K.
Models: We conduct our experiments using two base models from the Qwen-2.5 family: (1) Qwen-2.5-0.5B and (2) Qwen-2.5-1.5B. and two base models from the Llama-3.2 family: (1)Llama-3.2-1B-Instruct and (2) Llama-3.2-3B-Instruct.
@@
Training Hyper-parameters: For all baseline implementations, we use the VERL pipeline (sheng2024hybridflow). All baselines are trained with consistent hyperparameter settings to ensure fair comparison. The training batch size is set to 64 and the mini-batch size to 8 for both datasets. To ensure an equal number of training samples across baselines, we sample 5 responses per question prompt for each method. During response rollouts, the temperature is set to 0.6. The maximum prompt length is set to 512 for GSM8K and 256 for countdown, while the maximum response length is fixed at 1024 for both datasets. The learning rate is set to 1e61e-6. For GRPO, we set the KL divergence coefficient to $\beta=1e-3$. In the case of Filtered-ISFT+–, we apply a constant weighting of 0.5 to both positive and negative responses throughout training. Model checkpoints are saved every 10 steps for evaluation on the test set. All experiments on the GSM8K dataset are conducted using a single A100 GPU, while Countdown experiments are run using two H100 GPUs. For the GSM8K dataset, both models of the Qwen and Llama family are trained for $145$ global time steps, corresponding to $5$ epochs. For the Countdown dataset, Qwen-2.5-0.5B is trained for 600 global time steps and Qwen-2.5-1.5B &\& Llama-3.2-3B-Instruct are trained for $300$ global time steps.
@@ 6.2 Results and Discussion
As a sanity check, we also evaluated Filtered-ISFT–, which uses only negative samples during fine-tuning. While its performance on GSM8K was on par with other baselines, it significantly underperformed on Countdown. This suggests that negative-only supervision may be inadequate for more challenging tasks such as Countdown, where positive examples likely play a critical role in guiding learning and improving generalization.
@@
We further evaluated larger models from the Llama-3.2 Instruct family (1B and 3B) on GSM8K, Countdown, and MATH500, with results reported in Table 2. Experiments with Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct show that F-ISFT+- and GRPO perform similarly on GSM8K and CD8K. On MATH500, GRPO performs better on the 1B model, while F-ISFT+- achieves slightly higher performance on the 3B model. Although F-ISFT+ has comparatively lower performance on the 1B model, all baselines achieve similar performance across datasets on the 3B model.
@@
Length bias analysis - The figures in 1 present results on the Countdown test dataset at different evaluation steps during the post-training of the Qwen-2.5-1.5B model. The left panel shows the number of correct and incorrect responses, where we observe that the number of correct responses increases steadily while the number of incorrect responses decreases across all baselines. The right panel shows the average token response length for both correct and incorrect responses. For correct responses, we find a slight increase in the average response length for GRPO and Filtered-ISFT+-, while it remains relatively constant for Filtered-ISFT+. For incorrect responses, the average length decreases initially but then increases substantially for GRPO and Filtered-ISFT+-, whereas Filtered-ISFT+ maintains comparable response lengths for both correct and incorrect responses.
@@ 7 Discussion and Conclusion
In this work, we critically examined the structural assumptions underlying the popular LLM-MDP formulation and analyzed how they undermine RL applications to LLM post-training. Focusing on the widely adopted GRPO algorithm introduced by DeepSeek (guo2025deepseek), we showed that by representing states as nothing more than sequences of actions, applying only terminal rewards, and uniformly distributing credit, GRPO closely resembles filtered iterative supervised fine-tuning. Our empirical results further demonstrated that variants of filtered iterative SFT, leveraging both positive and negative samples, achieve performance comparable to GRPO. We also observed that RL-trained LLMs tend to generate longer responses, an effect often misattributed to improved reasoning capabilities. Our findings suggest that this phenomenon actually arises from biases in the training setup, specifically the uniform distribution of advantages across tokens combined with length-based scaling effects.
@@
A key implication of our analysis is that misunderstanding the consequences of these structural assumptions has led to a number of downstream research directions that attempt to correct for symptoms of the formulation, rather than addressing the root cause. For example, recent works have proposed explicitly optimizing for shorter reasoning traces by introducing length penalties in the reward function (arora2025training; team2025kimi) or by avoiding training on longer incorrect responses indirectly, e.g. by sorting the responses by length and correctness  (shrivastava2025sample). While these methods appear to mitigate the tendency of RL-trained LLMs to produce excessively long outputs, they do so without recognizing that such length biases are intrinsic to the degenerate reward allocation mechanism of the current LLM-MDP framework. As a result, these interventions risk obscuring the fundamental issue and may unintentionally hinder the development of more principled approaches.
A promising direction may be to revisit the MDP formulation itself. The 2-LLM framework described in (valmeekamsystematic) provides one such avenue. In this formulation, a chain-of-thought (CoT) LLM policy generates intermediate reasoning steps as actions conditioned on the problem prompt, while a main LLM generates the final solution conditioned on both the prompt and the CoT actions. This setup more faithfully reflects the sequential decision-making structure assumed in RL and enables the design of reward functions that meaningfully shape intermediate reasoning. By adopting such alternatives, the full expressive capacity of RL can be more effectively leveraged in guiding LLMs toward better reasoning strategies.
To conclude, our critique is directed at the degenerate LLM-MDP formulation rather than at RL itself. Under the current restrictive assumptions, the potential of RL cannot be fully realized. We hope that our analysis motivates exploration of alternative formulations that can better unlock the benefits of RL for LLM post-training.
Acknowledgment
@@
This research is supported in part by grants from ONR (N00014-25-1-2301 and N00014-23-1-2409), DARPA (HR00112520016), DoD RAI (via CMU subcontract 25-00306-SUB-000), an Amazon Research Award, and a generous gift from Qualcomm. We thank Professors Dimitri Bertsekas and Angelia Nedic for their helpful discussions.
@@ A.3 Length Bias - Experimental Results
These observations support our analysis from Sections A.1 and  A.2. Specifically, the average length of incorrect responses increases as training progresses, whereas the length of correct responses remains relatively stable. Moreover, the number of correct responses consistently rises, while the number of incorrect responses declined. Thus, we analyze the increase in average response length by separating it into contributions from correct and incorrect responses. Our findings show that the increase is mainly caused by the incorrect responses. This can be explained by the structural assumptions, particularly by uniform advantage distribution across tokens and scaling the advantage by length during training. Moreover, we find no evidence that longer responses lead to better reasoning and that the reasoning abilities could emerge due to increased response length

### v4

@@ 4.2 Decomposing by Positive and Negative Responses
Figure 1: Equivalence between Filtered Iterative Supervised Fine-Tuning (FISFT) and GRPO: Starting from supervised fine-tuning, introducing on-policy sampling, verifier-based filtering, and weighted updates leads to FISFT, which closely matches the update structure of DeepSeek-R1’s GRPO. Conversely, when RL is formulated with simplifying assumptions such as degenerate MDP states, uniform reward splitting, and group-based advantages, the GRPO update rule becomes equivalent to FISFT. The figure illustrates how both pipelines converge to the same underlying training dynamics from opposite directions.
@@
In this formulation, the terms $A_{q,i}^{+}$ and $A_{q,i}^{-}$ serve as adaptive weights for the positive and negative components of the objective, respectively. When rewards are binary, these weights which are computed from the mean and standard deviation of rewards within a group, become function of model’s probability of succeess in finding a correct solution. Let p, be the probability of succeess. Specifically, Mroueh (2025) shows that higher weight is assigned to correct responses when $p<0.5$ and to incorrect responses when $p>0.5$. with equal weighting at $p=0.5$. In contrast, FISFT assigns equal weights to correct and incorrect responses, in Equation  8.
@@
Fig. 1 provides an illustration that SFT with on-policy modifications, such as clipping, importance sampling, and filtering the groups when all the respones are correct or incorrect, produces updates similar to DeepSeek-R1’s GRPO, while GRPO itself, derived from RL on a degenerate MDP under structural assumptions such as state concatenation and uniform reward splitting, has updates close to FISFT.
@@ 5 Length Bias in GRPO
This length bias arises from structural assumptions in the learning framework, specifically, dividing the advantage uniformly across tokens and scaling it with response length. We empirically show that this bias also appears in F-ISFT+-, similar to GRPO. Notably, F-ISFT+-, which is trained on both correct and incorrect responses, produces longer outputs than F-ISFT+, which is trained only on positive samples. Yet, the longer responses are often interpreted as signs of ’learning to reason’ (Guo et al., 2025), even though they arise from disproportionately longer incorrect responses caused by structural assumptions in the degenerate LLM-MDP framework. We also provide further analysis on length bias in the Appendix A.
@@ 6.1 Experimental Setup
Datasets: We present our analysis on two datasets - 1) the GSM8K dataset (Cobbe et al., 2021), which is a widely used benchmark dataset consisting of grade school math problems, designed to evaluate the reasoning capabilities of large language models. It contains 8.5K problems, each paired with a question and an answer. The dataset is divided into 7.5K training problems and 1K test problems. And, 2) Countdown dataset (Countdown., ), which is a generalized version of the classic 24 Game (Yang et al., 2022), where the objective is to combine a set of input numbers using basic arithmetic operations (+, –, ×, ÷) to reach a specified target number. In this dataset, each problem consists of 3 to 4 two-digit input numbers, with the target number also being a two-digit number. The dataset contains 9K examples, split into 8K training instances and 1K test instances. We also report the accuracy on the MATH-500 dataset, a 500-problem test set evaluated using models trained on GSM8K.
Models: We conduct our experiments using base models from the Qwen: (1) Qwen-2.5-0.5B (2) Qwen-2.5-1.5B. and (3) Qwen3-8B and Instruct models from the Llama and Deepseek: (1)Llama-3.2-1B-Instruct (2) Llama-3.2-3B-Instruct. and (3) Deepseek-math-7B-Instruct
@@
The training hyper-parameters and complexity details are provided in the appendices B and C
@@ 6.2 Results and Discussion
Methods Deepseek-math-7B-Instruct Qwen3-8B
GSM8k GSM8k
Base Model $00.07$ $65.27$
GRPO $82.4$ $\mathbf{92.1}$
PPO $81.9$ $91.8$
Filtered-ISFT+- $\mathbf{83.7}$ $91.5$
Table 2: Performance of GRPO, PPO and Filtered-ISFT+- on test dataset of GSM8k for Deepseek-math-7B-Instruct and Qwen3-8B models. The highest score is highlighted in bold.
As a sanity check, we also evaluated Filtered-ISFT, which uses only negative samples during fine-tuning. While its performance on GSM8K was on par with other baselines, it significantly underperformed on Countdown. This suggests that negative-only supervision may be inadequate for more challenging tasks such as Countdown, where positive examples likely play a critical role in guiding learning and improving generalization.
Table 2 shows the performance of GRPO, PPO, and F-ISFT+- on the GSM8K dataset for the larger models DeepSeek-Math-7B-Instruct and Qwen3-8B. From the results, it is evident that F-ISFT+- performs comparably to GRPO even at these larger scales. This further supports our claim that the behavior we analyze is not limited to small models and persists across stronger base models as well. Additionally, both GRPO and F-ISFT+- achieve similar performance to PPO while being more scalable, as they do not require a value function.
@@
We further evaluated the performance of all baselines on additional language models - the Llama-3.2 Instruct family models (1B and 3B) on GSM8K, Countdown, and MATH500, with results reported in Table 3. Experiments with Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct show that F-ISFT+- and GRPO perform similarly on GSM8K and CD8K. On MATH500, GRPO performs better on the 1B model, while F-ISFT+- achieves slightly higher performance on the 3B model. Although F-ISFT+ has comparatively lower performance on the 1B model, all baselines achieve similar performance across datasets on the 3B model.
@@
Length bias analysis - The figures in 2 present results on the Countdown test dataset at different evaluation steps during the post-training of the Qwen-2.5-1.5B model. The left panel shows the number of correct and incorrect responses, where we observe that the number of correct responses increases steadily while the number of incorrect responses decreases across all baselines. The right panel shows the average token response length for both correct and incorrect responses. For correct responses, we find a slight increase in the average response length for GRPO and Filtered-ISFT+-, while it remains relatively constant for Filtered-ISFT+. For incorrect responses, the average length decreases initially but then increases substantially for GRPO and Filtered-ISFT+-, whereas Filtered-ISFT+ maintains comparable response lengths for both correct and incorrect responses.
@@ 7 Discussion and Conclusion
In this work, we critically examined the structural assumptions underlying the popular LLM-MDP formulation and analyzed how they undermine RL applications to LLM post-training. Focusing on the widely adopted GRPO algorithm introduced by DeepSeek (Guo et al., 2025), we showed that by representing states as nothing more than sequences of actions, applying only terminal rewards, and uniformly distributing credit, GRPO closely resembles filtered iterative supervised fine-tuning. Our empirical results further demonstrated that variants of filtered iterative SFT, leveraging both positive and negative samples, achieve performance comparable to GRPO. We also observed that RL-trained LLMs tend to generate longer responses, an effect often misattributed to improved reasoning capabilities. Our findings suggest that this phenomenon actually arises from biases in the training setup, specifically the uniform distribution of advantages across tokens combined with length-based scaling effects.
Utility of analysing GRPO from FISFT perspective:
@@
A key implication of our analysis is that misunderstanding the consequences of these structural assumptions has led to a number of downstream research directions that attempt to correct for symptoms of the formulation, rather than addressing the root cause. For example, recent works have proposed explicitly optimizing for shorter reasoning traces by introducing length penalties in the reward function (Arora and Zanette, 2025; Team et al., 2025) or by avoiding training on longer incorrect responses indirectly, e.g. by sorting the responses by length and correctness  (Shrivastava et al., 2025). While these methods appear to mitigate the tendency of RL-trained LLMs to produce excessively long outputs, they do so without recognizing that such length biases are intrinsic to the degenerate reward allocation mechanism of the current LLM-MDP framework.
While some works attempt to address sequence length issues, many variants of GRPO focus on modifying the advantage estimation Shao et al. (2025); Zhao et al. (2025). Our analysis shows that, under structural assumptions within degenerate MDP, GRPO is equivalent in performance to Filtered Iterative SFT. In this FISFT setting, explicit advantage estimates are not required, instead, a verifier signal is sufficient to distinguish correct and incorrect responses.
This perspective helps explain several recent empirical results. Spurious Rewards Shao et al. (2025) shows that random rewards can lead to strong performance in Qwen models. Additionally, Chen et al. (2025) shows that random rewards can facilitate exploration in stronger models. Our work clarifies these results by showing that such random rewards primarily affect advantage estimates that are not essential for performance. This is illustrated by FISFT, which achieves performance comparable to GRPO using equal weights for correct and incorrect responses. Unlike random rewards, whose effectiveness depends on the base model’s ability to generate correct responses, as shown in Chen et al. (2025), FISFT achieves comparable performance across model families and scales by assigning equal weight to correct and incorrect responses.
Similarly, GMPO (Zhao et al., 2025) replaces the group average mean in GRPO with a geometric mean and achieves comparable performance, which effectively corresponds to reweighting correct and incorrect responses without changing the underlying training dynamics.
Contemporary works have also examined training dynamics through ablations such as RAFT++ and Reinforce-Rej Xiong et al. (2025). RAFT++ trains only on positive samples and does not filter groups where all samples are correct, while Reinforce-Rej applies filtering but excludes negative samples; as a result, both slightly underperform GRPO. In contrast, FISFT+-, motivated by our systematic analysis, incorporates both positive and negative samples with filtering and achieves performance comparable to GRPO.
To conclude, our work shows the degeneracy in LLM MDP formulation and the structural assumptions that lead to the equivalence between GRPO and Filtered Iterative SFT. By analyzing GRPO and its variants through the lens of FISFT, our analysis provides a clear explanation for length increase and the training dynamics of GRPO.
8 Acknowledgements
@@
This research is supported in part by ONR grant N0001423-1-2409, DARPA grant HR00112520016, and gifts from Qualcomm, J.P. Morgan and Amazon. We thank the entire Yochan group for helpful discussions, and Prof. Dimitri Bertsekas and Prof. Angelia Nedic for their insights on reinforcement learning.
@@ A.3 Length Bias - Experimental Results
These observations support our analysis from Sections A.1 and  A.2. Specifically, the average length of incorrect responses increases as training progresses, whereas the length of correct responses remains relatively stable. Moreover, the number of correct responses consistently rises, while the number of incorrect responses declined. Thus, we analyze the increase in average response length by separating it into contributions from correct and incorrect responses. Our findings show that the increase is mainly caused by the incorrect responses. This can be explained by the structural assumptions, particularly by uniform advantage distribution across tokens and scaling the advantage by length during training. Moreover, we find no evidence that longer responses lead to better reasoning and that the reasoning abilities could emerge due to increased response length.
Appendix B Training Hyper-parameters
For all baseline implementations, we use the VERL pipeline [18]. All baselines are trained with consistent hyperparameter settings to ensure fair comparison. The training batch size is set to 64 and the mini-batch size to 8 for both datasets. To ensure an equal number of training samples across baselines, we sample 5 responses per question prompt for each method. During response rollouts, the temperature is set to 0.6. The maximum prompt length is set to 512 for GSM8K and 256 for countdown, while the maximum response length is fixed at 1024 for both datasets. The learning rate is set to 1e61e-6. For GRPO, we set the KL divergence coefficient to $\beta=1e-3$. In the case of Filtered-ISFT+–, we apply a constant weighting of 0.5 to both positive and negative responses throughout training. Model checkpoints are saved every 10 steps for evaluation on the test set. All experiments on the GSM8K dataset are conducted using a single A100 GPU, while Countdown experiments are run using two H100 GPUs. For the GSM8K dataset, both models of the Qwen and Llama family are trained for $145$ global time steps, corresponding to $5$ epochs. For the Countdown dataset, Qwen-2.5-0.5B is trained for 600 global time steps and Qwen-2.5-1.5B &\& Llama-3.2-3B-Instruct are trained for $300$ global time steps. For the larger language models (DeepSeek-Math-7B-Instruct and Qwen3-8B) trained on GSM8K, we keep all hyperparameters the same as detailed above, except that instead of full-parameter fine-tuning, we apply LoRA fine-tuning for both models with rank 16 and alpha 32. The training is performed on two H100 GPUs for both the larger model for a total of 200 global training steps.
Appendix C Implementation Complexity and Training Dynamics
The implementation complexity of Filtered-ISFT+-, in terms of both space and computation, is roughly the same as GRPO. Similar to GRPO, at each training iteration we sample a batch of question prompts from the dataset and generate n responses per question prompt from the current model policy and label each response as positive or negative based on the correctness using an external verifier, store them in a buffer, and perform multiple gradient updates on that buffer. The updates simply increase the log-likelihood of tokens in positive responses and decrease the log-likelihood of tokens in negative responses.
Further in implementation, similar to GRPO, F-ISFT variants do not consider groups in which all responses are correct or all responses are incorrect. These cases correspond to empirical estimates of the base policy’s probability of success being 1 and 0, respectively. [13] provides a detailed analysis showing that these cases p = 0 & 1, correspond to fixed-point iterations of the objective function and therefore do not change the policy. Consequently, we filtered out groups containing only correct or only incorrect responses. In addition, [22] performed ablation studies on these groups and observed a decrease in performance when included.
In terms of training dynamics, both methods achieve similar performance, but F-ISFT+- demonstrates greater stability compared to GRPO. As shown in Equation 8, F-ISFT+- assigns equal weight to correct and incorrect responses during training, whereas GRPO uses adaptive weights based on empirical estimates of the mean and variance under the given policy, which depends on sampling
FISFT effectively functions as an on-policy variant of SFT. With the additional inclusion of importance sampling for on-policy correction, clipping, and the filtering of groups where all responses are correct or incorrect. We make the comparison that, under structural assumptions, GRPO becomes equivalent to a modified form of SFT that iteratively trains on self-generated data, incorporating modifications such as clipping, importance sampling, and group filtering when all responses are either correct or incorrect.